Ex-H) I found out that skewness is defined as 1. For this pmf 3 57 x) 0.1 0.8 0.1 Compute the skewness(20 points). 2. For this pdf. 0 else Compute the skewness (20 points).
Answer #1

SOLUTIOU 3 0 H= E(X) Ci = 0.9 t 20 +4.9 25.8 = (33 XOI) + (S3x08) + (73x01) ニ(27 x 0-1) +(125 XO-8) + (343×0-1) 2门+ 100+343 1Now , 25.8 (5) 25. 8-2S 0. 8 3 0 8 x0.894427 131-12-125 0.8% 0.89 4427 skew ness o el 3 8x3 24 1-083332 32 32 3 .. [55-35] : 40 [3125-243] 110 = 2882 ะ 72 , 05 40 = 17-40833) = 17-16-6733 0.3267 72-OS-(3x4 og3 3x 0.3267 )-(4.0